Q8: Terry and Associates conducted a study of service times at the drive-up window of fast-food...

Q8:

Terry and Associates conducted a study of service times at the drive-up window of fast-food restaurants. The average service time at Popeye's restaurants was 2.78 minutes.

A: What is the probability that a customer's service time is less than 3 minutes?

B: What is the probability that a customer's service time is more than 6 minutes?

C: What is the probability that a customer's service time is more than 2.78 minutes?

Question 8 options:

.6601, .1155, .3679

.5130, .1655, .5

.5130, .1655, .3679

We don't have enough information to completely answer this question.

.487, .1655, .5
Answer #1

Ans:

A)

P(T<3)=1-exp(-3/2.78)=0.6601

B)

P(T>6)=exp(-6/2.78)=0.1155

C)

P(T>2.78)=exp(-2.78/2.78)=exp(-1)=0.3679

Correct option is:

.6601, .1155, .3679